Horizon Inc. has a Canadian receivables contract for $200,000 due in 270 days. The firm has been approached by a factoring firm that offers to purchase the receivables at a 12.6% per annum discount plus a 1.00% charge for a nonrecourse clause. What is the annualized percentage all-in-cost of this factoring alternative?
Answer #1

charge for nonrecourse charge=200000*1%=$2000

Present value of discounted cash flow @12.6%= (200000-2000)/(1+12.6%*270/360)=$180904.5 (considering 1 year=360 days)

Total cost =(200000-180904.5)/200000=9.55% for 270 days

So, Annualized percentage all-in cost= (1+9.55%/270)^360-1=13.577%